LASER SPECTROSCOPY OF YBr: ROTATIONAL ANALYSIS OF THE C^{1}Σ − X^{1}Σ SYSTEM \underline{J. W-H. LEUNG}, JINGHUA DAI and A. S-C. CHEUNG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.